# Data Scoping: Effectively Learning the Evolution of Generic Transport PDEs

### v1

@@ Abstract
Transport phenomena (e.g., fluid flows) are governed by time-dependent partial differential equations (PDEs) describing mass, momentum, and energy conservation, and are ubiquitous in many engineering applications. However, deep learning architectures are fundamentally incompatible with the simulation of these PDEs. This paper clearly articulates and then solves this incompatibility. The local-dependency of generic transport PDEs implies that it only involves local information to predict the physical properties at a location in the next time step. However, the deep learning architecture will inevitably increase the scope of information to make such predictions as the number of layers increases, which can cause sluggish convergence and compromise generalizability. This paper aims to solve this problem by proposing a distributed data scoping method with linear time complexity to strictly limit the scope of information to predict the local properties. The numerical experiments over multiple physics show that our data scoping method significantly accelerates training convergence and improves the generalizability of benchmark models on large-scale engineering simulations. Specifically, over the geometries not included in the training data for heat transferring simulation, it can increase the accuracy of Convolutional Neural Networks (CNNs) by 21.7 % and that of Fourier Neural Operators (FNOs) by 38.5 % on average.
Figure 1: The target problem. (a) The deep learning architecture inevitably expands the scope of input data used for the prediction at one position as the number of layers increases, which contradicts the local-dependency assumption for a generic transport system. (b) The data scoping method proposed in this paper can ensure that the scope of the input data stays constant regardless of the number of layers, which decouples the expressiveness and local dependency of neural networks.
@@ 1 Introduction
Generic transport equations, a set of time-dependent partial differential equations (PDEs), describe the evolutions of the extensive properties in a physical system, such as mass, momentum, and energy. Derived from the conservation laws governing these properties, these equations are fundamental to understanding various physical phenomena in science and engineering, such as mass diffusion equations, heat transfer equations, and Navier–Stokes equations. The high-fidelity and rapid simulations of complex physical systems based on generic transport equations play a critical role in addressing design and prediction challenges in diverse fields where the solutions of many instances of PDE in different coefficients, initial conditions (IC) and boundary conditions (BC) are required, including airplane design [21], metal additive manufacturing control [5], weather forecasting [25], drug delivery [3], and pandemic outbreak modeling [26]. Traditionally, solving these PDEs in discretized forms, such as finite difference methods, finite element methods, and finite volume methods, incurs a cubic relationship between domain resolution and computation cost [13], which means that a 10-fold increase in resolution leads to a thousandfold increase in the computational cost. This computational challenge becomes apparent when considering realistic problems. For instance, the chord length of an airplane is about 2 m while the smallest eddy scale is $O(10^{-6})$ m. The advancements in computation infrastructure, such as GPUs, TPUs, and the relevant parallel computing platform, have paved the way for the success of machine learning (ML). This, in turn, signals a paradigm shift in scientific computation, with ML techniques emerging as valuable tools for addressing the limitations posed by traditional numerical frameworks.
@@
This paper introduces a data scoping method to enhance the generalizability of data-driven models predicting time-dependent physics properties in generic transport problems by decoupling the expressiveness and local-dependency of the neural operator, as illustrated in Figure 1. Our major contributions can be summarized as follows.
  • We define the local-dependency of generic transport PDEs in the form of a neural operator. From that, the incompatibility between the current deep learning architecture and the local-dependency of generic transport PDEs is demonstrated.
@@
  • We establish a data scoping method that solves the incompatibility mentioned above. It serves as both a preprocessing step and a postprocessing step. As a preprocessing step, it transforms the domain into information windows before the model makes predictions, and as a postprocessing step, it integrates these information windows. The linear time complexity of the method underscores its efficiency and scalability across different problem sizes.
  • We apply our method to two prominent neural network types — CNN and FNO — operating on regular grids. Numerical experiments are done over the data generated from the mass diffusion, fluid mechanics, and heat transfer with various coefficients and geometries. Our results demonstrate a significant enhancement in the convergence rate and generalization capabilities of these models.
2 Background
Physics-informed neural networks
Physics-informed neural networks (PINNs) have demonstrated the ability to learn the smooth solutions of known nonlinear PDEs with little or no data by incorporating PDE residuals into the training loss [27]. This approach has proven particularly valuable in solving inverse problems, enabling the identification of unknown coefficients in governing equations [11]. However, it is important to note that a single PINN model is typically trained to learn one specific instance of a PDE with specific coefficients, IC and BC [14]. Consequently, retraining the PINN model for a new instance of PDEs is necessary, and the associated large training cost stands out as a major limitation of the PINN framework [10]. This limitation poses a challenge to the generalizability and efficiency of PINNs, hindering their ability to overcome computation bottlenecks posed by traditional numerical methods.
Data-driven models
Data-driven models, which learn general physical patterns solely from data, hold significant promise for overcoming computation bottlenecks in approximating the solutions to a family of PDEs with various coefficients, IC and BC. [39, 1, 17, 7, 31, 13, 34, 36]. Despite the abundance of these such models, they commonly share a deep learning architecture characterized by multiple layers of basic neural networks, including Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Convolutional Neural Networks (GCNNs), and Fourier Neural Operators (FNOs). The expressiveness and nonlinearity of these models are realized through iterative forward computations across multiple layers. The design of the basic neural networks is specifically tailored for compatibility with matrix operations, efficiently accelerated in parallel on GPUs or TPUs. These ML models exhibit remarkable computational cost reductions, often on the order of three magnitudes, compared to traditional numerical methods, once they are well-trained [8]. However, the performance of data-driven is intricately tied to the quality of the data they are trained on. Notably, the training data for these models is predominantly sourced from traditional numerical methods or experimental observations, both of which can be prohibitively expensive to obtain [28]. This cost factor becomes particularly pronounced when aiming for a model capable of generalization to various PDE coefficients, IC, and BC for realistic problems. While there have been notable efforts to enhance the generalizability of data-driven models by tailoring model structures[32, 20, 33, 35], the deep learning architecture applied in these models hinders their generalizability because it is incompatible with the local-dependency of generic transport PDEs as elaborated below.
Local-dependency of generic transport systems
A data-driven model for generic transport PDEs can be formulated as a neural operator, mapping the current status of the system into the status at the next time step. The evolution of the system can be obtained through iterative updates of the status using this neural operator. In this paper, we assume that the operator only takes the current system status to make the prediction. Note that there is another line of methods that takes multiple previous system statuses [17, 22, 38]. While it can improve prediction accuracy, it requires more data and increases the training burden, which causes scalability problems for large-scale simulations.
In the context of generic transport systems, the speed of information propagation is limited. This implies that the physics properties at a position in the next time step depend on the current status of its neighbors, which is called local-dependency. The size of the neighborhood is determined by the characteristic information speed and the time step. The neighborhood is called the information window, or simply “window” in this paper. So, the neural operator should also be local dependent. If the input scope is larger than the information window, it would incur noises that distract the model from capturing the true physics pattern in the data. While CNN and GCNN have the property of local-dependency, the deep learning architecture weakens such property because the scope of the input data used to make the prediction at a position is enlarged as the layer number increases, as illustrated in Figure 1 (a) and will be further discussed in Section 3.3.
Domain decomposition methods
@@ 2 Background
The concept of data scoping in this paper shares similarities with domain decomposition methods commonly used in traditional numerical approaches. Domain decomposition methods aim to alleviate the hardware demands of solving a large domain by partitioning it into smaller regions that can be solved in parallel with defined interface conditions [9, 30]. A series of machine learning models have incorporated the concept of domain decomposition [15, 16, 10]. These models decompose the domain into subdomains, each assigned a PINN model that is trained independently. Information exchange between subdomains is facilitated by adjusting the boundary term or incorporating interface conditions in the training loss. Although these methods enhance computation speed by solving subdomains in parallel, they share limitations inherent in PINNs—specifically, they are tailored to a particular instance of a PDE and involve a time-consuming training process. Our data scoping method also supports parallel computation while it is designed for data-driven models that approximate the solutions of a family of PDEs.
3 Local-dependent Neural Operator
@@
PDEs can be viewed as nonlinear operators that map between Banach function spaces. We formulate the data-driven models as the nonlinear operators approximating a family of PDEs following the work of Li et. al.[17].
3.1 Solution Operator of Generic Transport PDEs
Formulation
@@ 3.1 Solution Operator of Generic Transport PDEs
The domain of a generic transport system in $d$-dimensional space is denoted as $D\subset\mathbb{R}^{d}$ which is a bounded open set. Let $d_{u}$ be the dimension of the physical properties transported in the system. Let $d_{a}$ be the dimension of the constant properties of a specific instance of PDEs, such as coefficients and boundary geometric features. Let $\mathcal{A}=\mathcal{A}(D;\mathbb{R}^{d_{a}})$ and $\mathcal{U}=\mathcal{U}(D;\mathbb{R}^{d_{u}})$ be Banach spaces of functions that take values in $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$, respectively. The constant properties of the system are denoted as $a\in\mathcal{A}$. The status of the system at time $t$ is denoted as $u^{t}\in\mathcal{U}$. For the convenience of formulation, the time dimension is discretized uniformly. We have $t=0,1,2,...,T$ with fixed timestep $\Delta t$ and maximum $T$. The evolution of the system from $t$ to $t+1$ can then be represented by a nonlinear operator $G^{{\dagger}}:\mathcal{A}\times\mathcal{U}\rightarrow\mathcal{U}$ in the way that
@@
Given $a$ and the initial status of the system $u^{0}$, $u^{t}$ can be calculated by $G^{{\dagger}}$ in an iterative way for all $t=0,1,2,...$. So, $G^{{\dagger}}$ can be viewed as the solution operator of a family of generic transport PDEs characterized by $\mathcal{A}$.
@@
Dicretization
The functions $a_{j}$ and $u_{j}$ are discretized for the ease of computation in practice. According to the definition in [14], the neural operator is mesh-independent, which means that with the same $G_{\theta}$ can be used for different discretizations and $C(G_{\theta}(a_{j},u^{t}_{j}),u^{t+1}_{j})$ should not have significant changes over different discretizations. In this paper, we limit the discussion to the same discretization, so we relax the mesh-independent requirements. We generalized the definition of the neural operator to any data-driven models that can take $a_{j}$ and $u^{t}_{j}$, output $u^{t+1}_{j}$ over the fixed discretization.
@@ 3.2 Local-dependency
The physical information in a generic transport system travels at a limited speed.
@@
So, we can define the local-dependent operator for the generic transport system in Definition 2.
@@
3.3 Incompatibility between deep learning and local-dependency
Here we explain why the deep learning architecture of neural operators weakens the local-dependency. A neural operator consists of multiple layers where each layer is a linear operator followed by a non-linear activation. The universal approximation theorem states that such architecture can accurately approximate any nonlinear operator [6]. The deep learning architecture of the neural operator for generic transport problems is formulated in an iterative way that
@@ 3.3 Incompatibility between deep learning and local-dependency
where $k_{\phi}$ is a family of parameterized periodic functions. However, under a neural operator that consist of multiple layers of the local-dependent convolutions, the local-dependent region at $x$ is larger than $U(x,\delta)$. Specifically, the size of the expanded local-dependent region is positive proportional to the layer number, which is stated in Theorem 1 and proved in Appendix A.
@@
So, we are in such a dilemma. Under the deep learning architecture, to increase the expressiveness of neural network to account for the nonlinearity of the generic transport PDEs, we need to increase the layer number. But increasing the layer number results in expanding the scope of the information used to make the prediction, which might violate the local-dependency of time-dependent generic transport PDEs as defined in Definition 2.
3.4 Data scoping
Instead of limiting the scope of the linear operator to one layer, our idea is to limit the scope of input data directly. The data scoping method decomposes the data so that each operator only works on the segmentation $u^{t}|_{U(x,\delta)}$ defined in Equation 3. So, now we have
@@ 3.4 Data scoping
Under this formulation, the calculation of $u^{t+1}_{j}(x)$ only involves the information in $U(x,\delta)$ no matter the scope of the linear operator $K_{\phi}$. To realize the segmentation $u^{t}|_{U(x,\delta)}$ efficiently, we developed a data scoping method to partition the data into windows with prescribed sizes and integrate the predictions over the individual windows into the whole domain.
Given a domain discretized by a grid, as illustrated in Figure 2 (a), to predict the physical properties in the next timestep at one position (colored in black), it is assumed that the local region colored in grey contains sufficient information. So, instead of inputting the whole domain into the ML model, we should only input the relevant local region (colored in grey) to make the prediction (colored in orange). Our domain decomposition algorithm partitions the domain evenly into smaller windows and has the ML model make the predictions at the centers of the windows. The details of the domain decomposition and its reverse, window patching, algorithms are illustrated in Figure 4 and explained in Section B.2. While one decomposition of the domain only generates the prediction at the center of the windows, as shown in Figure 2 (b), we need decompositions over an expanded domain as illustrated in Figure 3 and detailed in Section B.1 to make the complete prediction as indicated in Figure 2 (c). The prediction integration algorithm illustrated in Figure 5 and detailed in Section B.3 explains how to obtain the prediction over the complete domain.

### v2

# Enforcing the Principle of Locality for Physical Simulations with Neural Operators

@@ Abstract
Time-dependent partial differential equations (PDEs) for classic physical systems are established based on the conservation of mass, momentum, and energy, which are ubiquitous in scientific and engineering applications. These PDEs are strictly local-dependent according to the principle of locality in physics, which means that the evolution at a point is only influenced by the neighborhood around it whose size is determined by the length of timestep multiplied with the speed of characteristic information traveling in the system. However, deep learning architecture cannot strictly enforce the local-dependency as it inevitably increases the scope of information to make local predictions as the number of layers increases. Under limited training data, the extra irrelevant information results in sluggish convergence and compromised generalizability. This paper aims to solve this problem by proposing a data decomposition method to strictly limit the scope of information for neural operators making local predictions, which is called data decomposition enforcing local-dependency (DDELD). The numerical experiments over multiple physical phenomena show that DDELD significantly accelerates training convergence and reduces test errors of benchmark models on large-scale engineering simulations.
@@ 1 Introduction
Figure 1: The target problem. (a) The deep learning architecture inevitably expands the scope of input data used for the prediction at one position as the number of layers increases, which is not compatible with the local-dependency assumption for a classical physics system. (b) DDELD method proposed in this paper can ensure that the scope of the input data stays constant regardless of the number of layers, which decouples the expressiveness and local dependency of neural networks.
@@
A broad variety of classical, time-dependent physical systems involve the evolution of mass, momentum, and energy. These relationships can be described by partial differential equations (PDEs) with local dependency under the assumption that the physical information travels at a limited speed through the physical medium. Solving these PDEs are fundamental to overcoming engineering challenges like airplane design [1, 2], additive manufacturing control [3], weather forecasting [4, 5], drug delivery [6], and pandemic outbreak modeling [7]. Traditional methods for solving these PDEs in discretized form include finite difference methods, finite element methods, and finite volume methods. However, all of these incurs a cubic relationship between domain resolution and computation cost [8], which means that a 10-fold increase in resolution leads to a 1000-fold increase for 3D problem in the computational cost. Advancements in computation infrastructure and parallel computing have paved the way for the success of machine learning (ML). This, in turn, signals a paradigm shift in scientific computation, with ML techniques emerging as valuable tools for addressing the computational limitations with sub-cubic costs [9, 10].
State-of-the-art methods
Physics-informed neural networks (PINNs) have demonstrated the ability to learn the smooth solutions of known nonlinear PDEs with little or no data by incorporating PDE residuals into the training loss [11]. This approach has proven particularly valuable in solving inverse problems, enabling the identification of unknown coefficients in governing equations [12]. However, it is important to note that a single PINN model is typically trained to learn one specific instance of a PDE with specific coefficients, initial conditions (IC) and boundary conditions (BC) [13]. Consequently, it is necessary to retrain the PINN model for every new PDE instance, and the associated large training cost stands out as a major limitation of the PINN framework [14]. This limitation poses a challenge to the generalizability and efficiency of PINNs, hindering their ability to overcome computational limitations posed by traditional numerical methods.
In addition, neural operators have emerged as an effective way to overcome computation bottlenecks in approximating the solutions for a family of PDEs by learning the mapping between function spaces from data, such as DeepONet [15] and Fourier Neural Operators (FNOs) [16]. Besides, conservative laws incorporated into neural operators have been found can improve learning process with limited data [17, 18, 19, 20, 21]. The expressiveness and nonlinearity of these models are realized through deep learning architecture, iterative forward computations across multiple linear operators. However, the multiple-layer architecture of deep learning applied in these models cannot focus on the local evolution patterns for time-dependent PDEs as elaborated below.
Unsolved problem
In the context of classical systems, information propagates at a limited speed. This implies that the physics properties at a position in the next time step depend on the current status of its neighbors, which is called local-dependency. An ML model for time-dependent PDEs can be formulated as a neural operator, mapping the current status of the system into the status at the next time step. So a neural operator approximating a family of time-dependent PDEs in classical systems should also have local-dependent property, which only utilizes a specified window of information around a position to make the local prediction. Even though some neural networks, such as CNN based methods [22, 23], Graph-CNN based methods [24, 25], and neural operator with localized kernels [26], have the property of local-dependency in one layer, the multiple-layer architecture will expand the size of information window unintentionally as the layer number increases as illustrated in Figure 1 (a) and will be further discussed in Section 3.2. Given a specific time step, the local information near the boundary of the enlarged window would be beyond its reach to have causal effects on the center of the window. Therefore, it only adds noise, which distracts the neural operator from capturing the true local physical patterns over the center of the window.
@@
This paper introduces a data decomposition method for neural operators, called data decomposition enforcing local-dependency (DDELD), to ensure strict local-dependency in making the predictions for time-dependent PDEs of classic physical systems, as illustrated in Figure 1. Our major contributions can be summarized as follows.
  • We demonstrate that the multiple-layer architecture is not suitable for local-dependency PDEs as demonstrated in Figure 1 (a).
@@
  • We establish a method to solve the incompatibility. It decomposes the domain into small windows based on local-dependency, and integrates these windows with linear time complexity, which underscores its efficiency and scalability across different problem sizes.
  • We apply DDELD to the benchmark neural operators over the data of complex fluid mechanics. Our results demonstrate a significant enhancement in the convergence rate and generalization capabilities of benchmark neural operators.
Remarks
@@
The method proposed in this paper shares similarities with domain decomposition methods commonly used in traditional numerical approaches. Domain decomposition methods aim to alleviate the hardware demands of solving a large domain by partitioning it into smaller regions that can be solved in parallel with defined interface conditions [27, 28, 29, 30]. A series of ML models have incorporated the concept of domain decomposition [31, 32, 14, 33]. These models decompose the domain into subdomains, each assigned a PINN model that is trained independently. Information exchange between subdomains is facilitated by adjusting the boundary term or incorporating interface conditions in the training loss. Although these methods enhance computation speed by solving subdomains in parallel, they share limitations inherent in PINNs—specifically, they are tailored to a particular instance of a PDE and involve a time-consuming training process. In contrast, DDELD proposed in this paper is designed for neural operators that approximate the solutions of a family of PDEs while also supports parallel computation.
2 Background of Neural Operators
@@
PDEs can be viewed as nonlinear operators that map between Banach function spaces. We formulate the ML models approximating a family of time-dependent PDEs as the nonlinear operators following the work of Li et. al.[16].
Neural operator formulation
@@
The domain of a classical physical system in $d$-dimensional space is denoted as $D\subset\mathbb{R}^{d}$ which is a bounded open set. Let $d_{u}$ be the dimension of the physical properties evolved in the system. Let $d_{a}$ be the dimension of the constant properties of a specific instance of PDEs, such as coefficients. Let $\mathcal{A}=\mathcal{A}(D;\mathbb{R}^{d_{a}})$ and $\mathcal{U}=\mathcal{U}(D;\mathbb{R}^{d_{u}})$ be Banach spaces of functions that take values in $\mathbb{R}^{d_{a}}$ and $\mathbb{R}^{d_{u}}$, respectively. The constant properties of the system are denoted as $a\in\mathcal{A}$. The status of the system at time $t$ is denoted as $u^{t}\in\mathcal{U}$. For the convenience of formulation, the time dimension is discretized uniformly. We have $t=0,1,2,...,T$ with fixed timestep $\Delta t$ and maximum $T$. The evolution of the system from $t$ to $t+1$ can then be represented by a nonlinear operator $G^{{\dagger}}:\mathcal{A}\times\mathcal{U}\rightarrow\mathcal{U}$ in the way that
@@
Given $a$ and the initial status of the system $u^{0}$, $u^{t}$ can be calculated by $G^{{\dagger}}$ in an iterative way for all $t=0,1,2,...$. So, $G^{{\dagger}}$ can be viewed as the solution operator of a family of time-dependent PDEs characterized by $\mathcal{A}$.
@@
Discretization
The functions $a_{j}$ and $u_{j}$ are discretized for the ease of computation in practice. According to the definition in [13], the neural operator is mesh-independent, which means that the same $G_{\theta}$ can be used for different discretizations and $C(G_{\theta}(a_{j},u^{t}_{j}),u^{t+1}_{j})$ should not have significant changes over different discretizations. In this paper, we limit the discussion to the same discretization, so we relax the mesh-independent requirements. We generalized the definition of the neural operator to any data-driven models that can take $a_{j}$ and $u^{t}_{j}$, output $u^{t+1}_{j}$ over the fixed discretization.
3 Incompatibility between Deep Learning and Local-dependency
@@ 3.1 Local-dependency
The physical information in a classical physical (non-quantum mechanics) system travels at a limited speed. In classical physics, the principle of locality states [34] that for a cause at one point to have an effect at another point, something in the space between those points must mediate the action. To exert an influence, something, such as a wave or particle, must travel through the space between the two points, carrying the influence. The special theory of relativity limits the maximum speed at which causal influence can travel to the speed of light. In a regular engineering system that involves the evolution of the distribution of mass, momentum, and energy, the physical information is usually driven by the effects of fluctuation, diffusion, and convection, whose speed is much slower than light. For example, in linear advection system, the information traveling speed can be characterized by the velocity field, and the information near the local up-winding region is preferred to take into account for evolution simulation. In this paper, we aims to develop a general method to all the time-dependent classical physical systems that the local-dependency can be strictly applied for ML models.
@@
So, we can define the local-dependent operator for the system in Definition 2.
@@
3.2 The more the layers, the weaker the local-dependency
Here we explain why the deep learning architecture of neural operators weakens local-dependency. A neural operator consists of multiple layers where each layer is a linear operator followed by a non-linear activation. The universal approximation theorem states that such architecture can accurately approximate any nonlinear operator [35]. The deep learning architecture of the neural operator for time-dependent PDEs can be formulated in an iterative way that
@@ 3.2 The more the layers, the weaker the local-dependency
where $k_{\phi}$ is a family of parameterized periodic functions. However, under a neural operator that consists of multiple layers of the local-dependent convolutions, the local-dependent region at $x$ is larger than $U(x,\delta)$. Specifically, the size of the expanded local-dependent region is positively proportional to the layer number, which is stated in Theorem 1 and is proved in Appendix A.
@@
This result presents the incompatibility. Under the deep learning architecture, to increase the expressiveness of neural network to account for the nonlinearity of the PDEs, we need to increase the layer number. However increasing the layer number results in expanding the scope of the information used to make the prediction, which might violate the local-dependency of time-dependent PDEs as defined in Definition 2.
4 Methodology
4.1 Method formulation
Instead of limiting the scope of the linear operator to one layer, we propose limiting the scope of input data directly. DDELD decomposes the data so that each operator only works on the segmentation $u^{t}|_{U(x,\delta)}$ defined in Equation 3. So, now we have
@@ 4.1 Method formulation
Under this formulation, the calculation of $u^{t+1}_{j}(x)$ only involves the information in $U(x,\delta)$ no matter the scope of the linear operator $K_{\phi}$. Note that the same $K_{\phi}$ is used for all segmentations. To realize the segmentation $u^{t}|_{U(x,\delta)}$ efficiently, we developed DDELD to partition the data into windows with prescribed sizes and integrate the predictions over the individual windows into the whole domain.
Given a domain discretized by a grid, as illustrated in Figure 2 (a), to predict the physical properties in the next timestep at one position (colored in black), it is assumed that the local region colored in grey contains sufficient information. So, instead of inputting the whole domain into the ML model, we should only input the relevant local region (colored in grey) to make the prediction (colored in orange). Our domain decomposition algorithm partitions the domain evenly into smaller windows and has the ML model make the predictions at the centers of the windows. The details of the domain decomposition and its reverse, window patching, algorithms are illustrated in Figure 4 and explained in Section 4.2. While one decomposition of the domain only generates the prediction at the center of the windows, as shown in Figure 2 (b), we need decompositions over an expanded domain as illustrated in Figure 3 and detailed in Section 4.3 to make the complete prediction as indicated in Figure 2 (c). The prediction integration algorithm illustrated in Figure 5 and detailed in Section 4.4 explains how to obtain the prediction over the complete domain.
The window size $\delta$ is selected as a hyper-parameter in the scale of a characteristic length $L_{c}$ for specific PDEs. It’s difficult to apply a generalized term regarding the choice of window size for all the PDEs, but we can still give suggestions on the choice of window size based on whether the PDE is convection or diffusion dominant, or if the derived physical property has a clear frequency character (for example, a peak in the energy spectrum). The details of window size determination are discussed in Section 4.5.